# Epistemic Diversity and Knowledge Collapse in Large Language Models

### v5

@@ 5 Measuring Epistemic Diversity
After obtaining $\mathcal{X}_{m{}_{t}}$, we must quantify cluster diversity to rank different models. We could simply count the number of clusters for each model, but that would discount the distribution of the clusters (e.g., for one model, there might be a long tail of clusters with only one claim in each cluster, which should rank lower compared to a model with fewer clusters that each contain many claims). Conversely, we could turn these counts into probabilities, obtaining a categorical distribution over clusters, and then calculate entropy Peterson (2025), but then the change in total number of clusters across topics would not scale properly (e.g., compared to another setting that has half as many clusters, the diversity metric would only lower by a maximum of one point).
We therefore use Hill diversity Hill (1973); Jost (2006) as our diversity index. Hill diversity offers a general way to measure the relative abundance of different categories of items in a sample, and satisfies the “replication principle” where relative changes in categories and abundances between samples result in proportionate changes to the score Roswell et al. (2021). Using the notation from § 3, we calculate the general Hill diversity as
$D(\mathcal{X}_{m_{t}})=\left(\sum_{i}^{\mathcal{X}_{m_{t}}}p_{i}(\frac{1}{p_{i}})^{l}\right)^{1/l},$
@@
Figure 3: Epistemic diversity vs. model release date. Each point is a single model, with lines connecting models of approximately the same size across released versions. Error bars are 95% boostrapped confidence intervals based on the HSD of each topic (N=155). Absolute diversity is low for all models compared to a very modest search baseline (top 20 Google search results for each topic). However, for all families except Qwen and most sizes, we see a trend of improved diversity.
@@ 6 Empirical Results
To what extent are LLMs exhibiting knowledge collapse? To answer, we pose several sub-questions focused on how epistemic diversity has changed over time (RQ1), how epistemic diversity is impacted by generation setting (RQ2), how size and model family impact epistemic diversity (RQ3), and how cultural context impacts both diversity and representation (RQ4).
@@ 6.1 Models, Topics, and Settings
We test each model with two different settings: (1) parametric memory with instruction-fine-tuning only (IFT) and (2) retrieval augmented generation (RAG). In addition, to compare the epistemic diversity of LLMs with that of a traditional web search, we include as a baseline 20 retrieved web pages from a Google search for each topic. [^8] Search performed on Aug 10, 2025, in the US region. We filter out content from social media, PDFs, and any page with fewer than 1000 characters. Furthermore, we use the search data as our RAG database by splitting each web page into paragraphs. For retrieval, we match prompts to the top paragraphs using S-BERT embeddings, including up to 1000 tokens of context from these retrieved paragraphs. Since many prompts match with similar top paragraphs in the search results we additionally shuffle the ranks of all paragraphs with a cosine similarity to the query above 0.35 to encourage different prompts to use different contexts. [^9] We calculated this as the mean cosine similarity across all paragraphs and prompts in the data. This way, we can observe how varying context impacts output diversity in the RAG setting. For generation, we use top $p$ sampling ($p=0.9$, temperature =1.0=1.0) and generate a maximum of 2,100 tokens per prompt ($\approx$ the median number of tokens on the web pages in our search results).
For rarefaction, we calculate the coverage of each sample for each setting on each topic using Equation 2, and rarefy all samples to the minimum. [^10] We found that by sampling 200 responses per LLM per topic, the coverage tends to vary widely between models (20%-80%), while for search it tends to be << 20%. Due to the low coverage achieved by the search baseline, we only rarefy search if the coverage is above the minimum achieved by an LLM. Therefore, all results with search should be considered as a lower-bound baseline; the true diversity of search relative to the LLMs tested will be higher.
@@
As our study is broad (27 models, two settings, and 200 prompt templates per model per topic), we constrain the number of selected topics for computational feasibility. We use a sample of 30 general topics from IssueBench Röttger et al. (2025) and 125 additional hand-curated topics. These include 10-13 country-specific topics about important figures and historical events to help measure the impact of cultural context on diversity. To ensure that these are well documented and known topics, we only select topics where their associated English Wikipedia page has a content rating of at least “C”. [^11] https://en.wikipedia.org/wiki/Wikipedia:Content_assessment Following IssueBench, these include a mix of both controversial and non-controversial topics, e.g., general topics for which there may be contentious views, like nuclear weapons and pornography, and events which may be subject to diverse and varied interpretations. All events occurred before the earliest knowledge cutoff of any model in our study. In total, we have 155 topics (see Appendix § A.3 for the list), yielding 1.7M responses prior to decomposition and 69.5M claims after decomposition. While the topics were hand-picked based on the above criteria, the selection process could nevertheless impact the generalizability of our results. However, if we can identify knowledge collapse even for this limited set of topics, this would already be cause for concern.
@@ 6.2 Epistemic Diversity Across Time
First, we focus on RQ1: How has epistemic diversity in LLMs changed over time? We measure epistemic diversity across the 155 topics and 27 models mentioned previously, and plot Hill-Shannon diversity (HSD) vs. model release date in Fig. 3. We plot traditional search diversity as a baseline and separately plot IFT and RAG to see the impact of generation setting. We connect models over time based on their relative size and version.
@@
That being said, when looking at epistemic diversity compared to the search baseline, LLM outputs have low diversity. One can generally expect to encounter more diverse information about these topics by reading through the top-20 web pages rather than prompting an LLM in different ways. The search baseline is, in fact, a quite weak baseline, as we use only the top 20 Google search results for each topic, compared to sampling from each language model with 200 unique prompts per topic. Further, as mentioned in § 6.1, the search baseline is under-estimated relative to the LLM results. Therefore, while it is encouraging that LLMs have become more epistemically diverse over time, their diversity relative to search baselines is often low, which may risk knowledge collapse.
@@
Intercept (IFT) 1054.77($\pm 242.65$) $p\ll 1\mathrm{e}{-3}***$
RAG +739.186($\pm 34.06$) $p\ll 1\mathrm{e}{-3}***$
Search +1311.14($\pm 1284.00$) p<0.05p<0.05*
Table 2: Estimated coefficients ($\hat{\beta}$) and $p$-values of a linear mixed effects model with the setting (IFT, RAG, Search) as fixed effects and the model as random effects. The dependent variable is HSD. Both RAG and the search baseline lead to statistically significantly more diverse responses than IFT.
Predictor $\widehat{\beta}$ $p$-value
Intercept (Med.) 1034.47($\pm 249.49$) $p\ll 1\mathrm{e}{-3}***$
Small +239.190($\pm 48.21$) $p\ll 1\mathrm{e}{-3}***$
Large -228.47($\pm 49.5$) $p\ll 1\mathrm{e}{-3}***$
Table 3: Estimated coefficients ($\hat{\beta}$) and $p$-values of a linear mixed effects model with the model size (Small, Medium, and Large) and setting (IFT and RAG) as fixed effects and the model/version as a random effect. The dependent variable is HSD. We find an inverse relationship between model size and diversity..
6.3 Epistemic Diversity Across Settings
Our next focus is on RQ2: What is the impact of generation setting on epistemic diversity? We look at epistemic diversity across three settings: parametric knowledge in instruction fine-tuned models (IFT), the same models with retrieval-augmented generation (RAG), and the traditional search engine baseline. From Fig. 3, we see that, on our topics, RAG appears to have a strong positive impact on diversity, and search tends to be better than both IFT and RAG. To test if this is statistically significant, we use a linear mixed effects regression model with HSD as the dependent variable, generation setting as a categorical fixed effect, and the model as a random effect in order to control for the baseline diversity of each model.
From Table 2, we see that both RAG and the traditional search baseline yield significantly more diverse outputs than relying on parametric memory with IFT, at least for our topics. This highlights the potential of RAG in ensuring that models are epistemically diverse going forward. However, this may depend on RAG databases remaining human written. If traditional search platforms and RAG knowledge bases become dominated by LLM-generated content, the diversity benefits of RAG could be erased, risking knowledge collapse. Therefore, given the benefits that RAG can endow to epistemic diversity, we recommend RAG databases remain diverse and additionally prevent contamination from an overabundance of LLM generated text.
@@ 6.4 Epistemic Diversity Across Models
Firstly, from Fig. 3, we see that model size appears to have an unintuitive negative impact on epistemic diversity. To determine if the effect is statistically significant, we use a linear mixed effects model with the HSD as the dependent variable, the model size binned into three categories (Small, Medium, and Large; see Table 4 for how these are determined) as fixed effects, the generation setting (IFT or RAG) as additional fixed effects, and the model version as random effects. The results in Table 3, highlight that the observed effect w.r.t. model size is statistically significant. This echoes contemporary work which has shown that larger models tend to memorize more Morris et al. (2025), can show greater bias Bai et al. (2025b), and can be less diverse Zhang et al. (2025) than smaller models. Therefore, in settings where epistemic diversity is important (e.g., in viewpoint/opinion generation), smaller models may be a better choice.
But how do different models compare? We measure model similarity by measuring the Jensen-Shannon divergence (JSD) between the distributions of meaning classes $p_{i}$ generated by the most diverse model from each family and average this across topics (Fig. 4). We observe that the overall divergence is high across models for our topics. This suggests that prompting multiple models may improve epistemic diversity. Additionally, we observe that the open-weight models are more similar to each other than to GPT5 or the search baseline. Finally, we see that the greatest divergence across models is with the search baseline, suggesting that traditional search and LLMs produce largely different distributions of claims, at least for our topics.
@@ 6.5 Epistemic Diversity Across Countries
Second, to what extent are English and local language knowledge represented by the claims generated by LLMs? To answer this, we match LLM claims to Wikipedia claims in both English and the country’s local language using InfoGap (Samir et al. 2024, see Appendix § A.5 for details). We restrict this experiment to 9 countries in our study where the local language Wikipedias are in the top 20 in terms of active users, with the exception of Saudi Arabia for which the local language pages had minimal content. In order to quantify the extent to which Wikipedia claims appear in model outputs, we measure the HSD of each model after filtering out claims that are not matched to either the English or local language Wikipedia. This is equivalent to “minimal representativeness” outlined by Peterson (2025), which defines that, for a set of items $i$ that are relevant to a particular task (in this case, claims which appear on either English or local language Wikipedia pages), “any item should have at least some chance of appearing in the LLM output.” Results are plotted as a bar chart in Fig. 6, aggregated across models and topics for each country.
For the selected topics, we observe a statistically significant knowledge gap between English and local languages for 5 out of 8 countries, while for the remaining 3, the gap is not statistically meaningful. In no case is local representation statistically significantly greater than English representation. Additionally, representation for the USA specific topics are statistically significantly greater than every other country. This suggests that current English LLMs may fail to present knowledge which has not been presented in English for certain countries, potentiating a risk in terms of knowledge erasure.
@@ 7 Discussion and Conclusion
Our work is the first broad study of knowledge collapse in LLMs, where we propose a new methodology for measuring epistemic diversity. We use this to conduct an empirical study of 27 LLMs across 155 topics, with 200 prompt variations.
@@
In terms of topical coverage, we use a combination of randomly selected and hand curated topics. We attempted to select topics that broadly cover general concepts, historical events, and important figures across 12 countries, including both controversial and non-controversial topics. However, manual selection of topics may inevitably impact the generalizability of our results. Additionally, for some countries in the study, selection of topics is performed from an outsider perspective. An improvement to our setup could be to cover more topics by identifying what people typically search for using LLMs and randomly sampling from that pool of queries. Other notions of popularity can also be examined in terms of Wikipedia page views or trending search keywords. Regarding our decomposition and clustering algorithms, while we find through manual and automated evaluation that the quality tends to be high, performance is not perfect. Therefore, there is some irreducible noise in the results. Next, we choose a RAG setup which is intended to simulate a real-world RAG scenario, while actual RAG implementations will inevitably vary. It would therefore be useful to examine the utility of RAG in a more systematic way, benchmarking several implementations. Finally, while epistemic diversity is important, it should also be contextualized with other important aspects of knowledge such as factuality and relevance.

### v6

@@ 5 Measuring Epistemic Diversity
After obtaining $\mathcal{X}_{m{}_{t}}$, we must quantify its diversity to rank different models. There are a number of possible choices to measure diversity based on categorical distributions Estève et al. (2025). We could simply measure richness, i.e., count the number of meaning classes generated by each model, but that would ignore how meaning classes are distributed (e.g., for one model, there might be a long tail of clusters with only one claim in each cluster, which should rank lower compared to a model with fewer clusters that each contain many claims). We could instead calculate entropy as in Peterson (2025), which is done as follows:
$H(\mathcal{X}_{m_{t}})=-\sum_{i}^{{|\mathcal{X}_{m_{t}}}|}p_{i}\log p_{i}\,$
Entropy takes into account how claims are distributed, since $H$ increases when there are more meaning classes that are more evenly distributed in $\mathcal{X}_{m_{t}}$. However, $H$ scales logarithmically, which is unintuitive when comparing different settings for how diverse their claims are (e.g., when the number of semantically equivalent meaning classes in a given setting is cut in half, $H$ can only lower by a maximum of one point in base 2).
An appropriate measure of diversity that accounts for both of these issues is Hill diversity Hill (1973); Jost (2006), which generalizes many diversity measures including entropy. It does so by offering a general way to measure the relative abundance of different categories of items in a sample, and satisfies the “replication principle” where relative changes in categories and abundances between samples result in proportionate changes to the score Roswell et al. (2021). Using the notation from § 3, we calculate the general Hill diversity as
$D(\mathcal{X}_{m_{t}})=\left(\sum_{i}^{|\mathcal{X}_{m_{t}}|}p_{i}(\frac{1}{p_{i}})^{l}\right)^{1/l},$
@@
This measure has an intuitive interpretation Roswell et al. (2021): first, from Equation 1 we can see that, since $e^{x}$ is a monotonically increasing function of $x$, HSD is maximized when the entropy is maximized, i.e., under a uniform distribution. Consider a distribution $\mathcal{X}_{m_{t}}$ generated by a model $m$ for topic $t$ where $|\mathcal{X}_{m_{t}}|=K$ (i.e., $K$ meaning clases). Under a uniform distribution, the HSD resolves to:
$D_{S}{}(\mathcal{X}_{m_{t}})$ $=\exp\{-\sum_{i}^{K}\frac{1}{K}\log\frac{1}{K}\}$
$=\exp\{\log K\}=K$
In other words, if $|\mathcal{X}_{m_{t}}|=K$ then $D_{S}{}(\mathcal{X}_{m_{t}})=\hat{K}\leq K$, with $D_{S}{}(\mathcal{X}_{m_{t}})$ maximized under a uniform distribution. When $\hat{K}<K$, we can then view the effective diversity of the distribution of claims generated by $m$ as equivalent to a new distribution $\mathcal{X}_{\hat{m}_{t}}$ generated by a different model $\hat{m}$ which has a uniform probability of generating one of $\hat{K}$ meaning classes. This implies that one can, for example, claim a model which generates a claim distribution with an HSD of $2\hat{K}$ is twice as diverse a model which generates a claim distribution with an HSD of $\hat{K}$. In the context of epistemic diversity, one can say that a model with an HSD of $\hat{K}$ is effectively as diverse as one which produces $\hat{K}$ claims with equal probability.
@@
Figure 3: Epistemic diversity vs. model release date. Each point is a single model, with lines connecting models of approximately the same size across released versions. Error bars are 95% boostrapped confidence intervals based on the HSD of each topic (N=155). Absolute diversity is low for all models compared to a very modest search baseline (top 20/40 Google search results for each topic). However, for all families except Qwen and most sizes, we see a trend of improved diversity.
@@ 6 Empirical Results
What is the extent of the risk of knowledge collapse with LLMs? To answer, we pose several sub-questions focused on how epistemic diversity has changed over time (RQ1), how epistemic diversity is impacted by generation setting (RQ2), how size and model family impact epistemic diversity (RQ3), and how cultural context impacts both diversity and representation (RQ4).
@@ 6.1 Models, Topics, and Settings
We test each model with two different settings: (1) parametric memory with instruction-fine-tuning only (IFT) and (2) retrieval augmented generation (RAG). In addition, to compare the epistemic diversity of LLMs with that of a traditional web search, we include as a baseline the web pages retrieved from Google search for each topic. This includes a baseline with 20 retrieved pages (Search-20) and one with an additional 20 pages, for a total of 40 (Search-40) per topic. [^9] Search-20 performed on Aug 10, 2025 and additional Search-40 results retrieved on January 1, 2026, both in the US region. We filter out content from social media, PDFs, and any page with fewer than 1000 characters. Furthermore, we use the Search-20 data as our RAG database by splitting each web page into paragraphs. For retrieval, we match prompts to the top paragraphs using S-BERT embeddings, including up to 1000 tokens of context from these retrieved paragraphs. Since many prompts match with similar top paragraphs in the search results we additionally shuffle the ranks of all paragraphs with a cosine similarity to the query above 0.35 to encourage different prompts to use different contexts. [^10] We calculated this as the mean cosine similarity across all paragraphs and prompts in the data. This way, we can observe how varying context impacts output diversity in the RAG setting. For generation, we use top $p$ sampling ($p=0.9$, temperature =1.0=1.0) and generate a maximum of 2,100 tokens per prompt ($\approx$ the median number of tokens on the web pages in our search results).
For rarefaction, we calculate the coverage of each sample for each setting on each topic using Equation 2, and rarefy all samples to the minimum. [^11] We found that by sampling 200 responses per LLM per topic, the coverage tends to vary widely between models (20%-80%), while for search it tends to be << 20%. Due to the low coverage achieved by both search baselines, we only rarefy search if the coverage is above the minimum achieved by any LLM. Therefore, all results with search should be considered as a lower-bound baseline; the true diversity of search relative to the LLMs tested will be higher.
@@
As our study is broad (27 models, two settings, and 200 prompt templates per model per topic), we constrain the number of selected topics for computational feasibility. We use a sample of 30 general topics from IssueBench Röttger et al. (2025) and 125 additional hand-curated topics. These include 10-13 country-specific topics about important figures and historical events to help measure the impact of cultural context on diversity. To ensure that these are well documented and known topics, we only select topics where their associated English Wikipedia page has a content rating of at least “C”, meaning that the article is substantial enough to at least be useful to the casual reader. [^12] https://en.wikipedia.org/wiki/Wikipedia:Content_assessment Following IssueBench, these include a mix of both controversial and non-controversial topics, e.g., general topics for which there may be contentious views, like nuclear weapons and pornography, and events which may be subject to diverse and varied interpretations. All events occurred before the earliest knowledge cutoff of any model in our study. In total, we have 155 topics (see Appendix § A.3 for the list), yielding 1.7M responses prior to decomposition and 69.5M claims after decomposition. While the topics were hand-picked based on the above criteria, the selection process could nevertheless impact the generalizability of our results. However, if we can identify knowledge collapse even for this limited set of topics, this would already be cause for concern.
@@ 6.2 Epistemic Diversity Across Time
First, we focus on RQ1: How has epistemic diversity in LLMs changed over time? We measure epistemic diversity across the 155 topics and 27 models mentioned previously, and plot Hill-Shannon diversity (HSD) vs. model release date in Fig. 3. We plot traditional search diversity as a baseline (Search-20 and Search-40) and separately plot IFT and RAG to see the impact of generation setting. We connect models over time based on their relative size and version.
@@
That being said, when looking at epistemic diversity compared to the search baseline, LLM outputs have low diversity. Only the best RAG models and GPT-5 are able to achieve comparable diversity to the (underestimated) Search-20 baseline; no model is as diverse as the stronger (but still underestimated) Search-40 baseline. This means that one can generally expect to encounter more diverse information about our topics by reading through the top web pages rather than prompting an LLM in different ways. The search baselines are also quite weak baselines, as we use only up to the top 40 Google search results for each topic based on simply using the topic as a search term (e.g., “democracy”). This is compared to sampling from each language model with 200 unique prompts per topic. Further, as mentioned in § 6.1, the search baselines are under-estimated relative to the LLM results. In absolute terms, we can thus conclude that the most diverse model in our study (GPT-5) is at least 18.6% less diverse than a traditional search in terms of the claims it generates (HSD of 3110 vs. 2621), at least for our topics. This is concerning given other recent work comparing knowledge acquisition with LLMs vs. search, which shows that reliance on LLMs leads people to have a shallower and less original understanding of topics compared to traditional search Melumad and Yun (2025). Therefore, while it is encouraging that LLMs have become more epistemically diverse over time, their diversity relative to search is still low, which may risk knowledge collapse if such LLMs come to replace other means of knowledge acquisition, like search, for most people. [^13] We compare these results to the results we would achieve using semantic similarity only with the approach from Artificial Hivemind Jiang et al. (2025) in Fig. 9, see Appendix § A.4.2 for experimental details.
6.3 Epistemic Diversity Across Settings
Our next focus is on RQ2: What is the impact of generation setting on epistemic diversity? We look at epistemic diversity across three settings: parametric knowledge in instruction fine-tuned models (IFT), the same models with retrieval-augmented generation (RAG), and the traditional search engine baselines. From Fig. 3, we see that, on our topics, RAG appears to have a strong positive impact on diversity, and search tends to be better than both IFT and RAG. To test if this is statistically significant, we use a linear mixed effects regression model with HSD as the dependent variable, generation setting as a categorical fixed effect, and the model as a random effect in order to control for the baseline diversity of each model.
@@ 6.3 Epistemic Diversity Across Settings
Intercept (IFT) 939.90($\pm 223.67$) $p\ll 1\mathrm{e}{-3}$
RAG +656.50($\pm 33.83$) $p\ll 1\mathrm{e}{-3}$
Search-20 +1392.83($\pm 1183.75$) $p<0.05$
Search-40 +2170.85($\pm 1183.57$) $p\ll 1\mathrm{e}{-3}$
Table 2: Estimated coefficients ($\hat{\beta}$) and $p$-values of a linear mixed effects model with the setting (IFT, RAG, Search-20, and Search-40) as fixed effects and the model as random effects. The dependent variable is HSD. RAG and both search baselines lead to statistically significantly more diverse responses than IFT.
From Table 2, we see that both RAG and search yield significantly more diverse outputs than relying on parametric memory with IFT, at least for our topics. This highlights the potential of RAG in ensuring that models are epistemically diverse going forward. However, this may depend on RAG databases remaining human written. If traditional search platforms and RAG knowledge bases become dominated by LLM-generated content, the diversity benefits of RAG could be erased, risking knowledge collapse. Therefore, given the benefits that RAG can endow to epistemic diversity, we recommend RAG databases remain diverse and additionally prevent contamination from an overabundance of LLM generated text.
@@ 6.4 Epistemic Diversity Across Models
Firstly, from Fig. 3, we see that model size appears to have an unintuitive negative impact on epistemic diversity. To determine if the effect is statistically significant, we use a linear mixed effects model with the HSD as the dependent variable, the model size binned into three categories (Small, Medium, and Large; see Table 4 for how these are determined) as fixed effects, the generation setting (IFT or RAG) as additional fixed effects, and the model version as random effects. The results in Table 3, highlight that the observed effect w.r.t. model size is statistically significant. This echoes contemporary work which has shown that larger models tend to memorize more Morris et al. (2025), can show greater bias Bai et al. (2025b), and can be less diverse Zhang et al. (2025); Rassin et al. (2024) than smaller models. We hypothesize that this occurs since larger models have greater capacity and thus a greater propensity to memorize content Morris et al. (2025), therefore narrowing the model’s output token distribution given prompts about the same topic and thus the space of possible responses to contain fewer, more likely claims. Therefore, in settings where epistemic diversity is important (e.g., in viewpoint/opinion generation which should rely on diverse supporting facts), smaller models may be a better choice.
Predictor $\widehat{\beta}$ $p$-value
Intercept (Med.) 910.34($\pm 228.07$) $p\ll 1\mathrm{e}{-3}$
Small +219.39($\pm 44.28$) $p\ll 1\mathrm{e}{-3}$
Large -185.93($\pm 45.46$) $p\ll 1\mathrm{e}{-3}$
Table 3: Estimated coefficients ($\hat{\beta}$) and $p$-values of a linear mixed effects model with the model size (Small, Medium, and Large) and setting (IFT and RAG) as fixed effects and the model/version as a random effect. The dependent variable is HSD. We find an inverse relationship between model size and diversity.
But how do different models compare? We measure model similarity by measuring the Jensen-Shannon divergence (JSD) between the distributions of meaning classes $p_{i}$ generated by the most diverse model from each family and average this across topics (Fig. 4). We observe that the overall divergence is high across models for our topics. This suggests that prompting multiple models may improve epistemic diversity. Additionally, we observe that the open-weight models are slightly more similar to each other than to GPT-5 and the search baselines. Finally, we see that the greatest divergence across models is with the search baselines, suggesting that traditional search and LLMs produce largely different distributions of claims, at least for our topics.
@@ 6.5 Epistemic Diversity Across Countries
Second, to what extent are English and local language knowledge represented by the claims generated by LLMs? To answer this, we match LLM claims to Wikipedia claims in both English and the country’s local language using InfoGap, a recent approach which uses GPT-5 and multilingual sentence embeddings to match claims between Wikipedia pages in different languages (Samir et al. 2024, see Appendix § A.6 for details). We restrict this experiment to 9 countries in our study where the local language Wikipedias are in the top 20 in terms of active users, with the exception of Saudi Arabia for which the local language pages had minimal content. We use only the topics which are relevant for those countries. In order to quantify the extent to which Wikipedia claims appear in model outputs, we measure the HSD of each model after filtering out claims that are not matched to either the English or local language Wikipedia. This is equivalent to “minimal representativeness” outlined by Peterson (2025), which defines that, for a set of items $i$ that are relevant to a particular task (in this case, claims which appear on either English or local language Wikipedia pages), “any item should have at least some chance of appearing in the LLM output.” Results are plotted as a bar chart in Fig. 6, aggregated across models and topics for each country.
For the selected topics, we observe a statistically significant knowledge gap between English and local languages for 5 out of 8 countries, while for the remaining 3, the gap is not statistically meaningful. In no case is local representation statistically significantly greater than English representation. Additionally, representation for the USA specific topics are statistically significantly greater than every other country. This suggests that current English LLMs may fail to present knowledge which has not been presented in English for certain countries, potentiating a risk in terms of knowledge erasure, and thus collapse.
@@ 7 Discussion and Conclusion
Our work is the first empirical study of the risk of knowledge collapse in LLMs, where we propose a new methodology for measuring epistemic diversity. We use this to conduct an empirical study of 27 LLMs across 155 topics, with 200 prompt variations.
@@
In terms of topical coverage, we use a combination of randomly selected and hand curated topics. We attempted to select topics that broadly cover general concepts, historical events, and important figures across 12 countries, including both controversial and non-controversial topics. However, manual selection of topics may inevitably impact the generalizability of our results. Additionally, for some countries in the study, selection of topics is performed from an outsider perspective. An improvement to our setup could be to cover more topics by identifying what people typically search for using LLMs and randomly sampling from that pool of queries. Other notions of popularity can also be examined in terms of Wikipedia page views or trending search keywords. We also use only 200 different prompt templates due to computational contstraints. Geater coverage could be achieved through more sampling.
Regarding our decomposition and clustering algorithms, while we find through manual and automated evaluation that the quality tends to be high, performance is not perfect. Therefore, there is some irreducible noise in the results. Additionally, we adopt a strict notion of semantic equivalence for clustering in our methodology, while many real world claims may have some partial overlap in information which could be useful to capture when measuring diversity. Stronger NLI approaches also exist, such as using an LLM, but we choose a validated and more efficient BERT based approach at the cost of some accuracy. To illustrate the cost, we perform $\approx$744M NLI inferences in our study; assuming $\approx$40 tokens per inference, with GPT-5, this would cost $\approx$\$44,640.
Next, we choose a RAG setup which is intended to simulate a real-world RAG scenario, while actual RAG implementations will inevitably vary. It would therefore be useful to examine the utility of RAG in a more systematic way, benchmarking several implementations. Finally, while epistemic diversity is important, it should also be contextualized with other important aspects of knowledge such as factuality and relevance.